OpenStudy (anonymous):

solve the equation 3.7x - 5.0 = -2.2x + 0.5

OpenStudy (anonymous):

X = 1

OpenStudy (anonymous):

5.9x=5.5 x=5.5/5.9

OpenStudy (anonymous):

oops lol sorry

OpenStudy (anonymous):

wait but isn't it 5.9/5.5?

OpenStudy (anonymous):

3.7x - 5.0 = -2.2x + 0.5 +2.2x +2.2x 5.9x-5.0=.5 +5.0 +5.0 5.9x=5.5 /5.9 /5.9 x = .93

OpenStudy (anonymous):

Thank you guys.

OpenStudy (anonymous):

I forgot to say round to the nearest hundredths, my bad.

OpenStudy (anonymous):

3.7x-5.0=-2.2x+0.5 3.7x+2.2x=5.0+0.5 5.9x=5.5 x=5.5/5.9 x=0.93 x=1